# Roadside Monocular 3D Detection Prompted by 2D Detection

### v3

@@ 4.4 Societal Impacts and Limitations
Societal impacts. While our method achieves significantly better performance than prior works on the benchmark datasets of roadside 3D object detection, we note potential risks if directly using our method for vehicle-vehicle communication and vehicle-infrastructure cooperative perception. Real-world applications should treat our approach as a module and evaluate the whole system for ultimate goals. Methods developed with existing datasets might be biased towards specific cities and countries. For example, in some countries, vehicles are running on the different side of the road from what are presented in the datasets. These are potentially negative societal impacts.
Limitations and future work. Our Pro3D exploits a scene prior; if camera pose is changed (e.g., due to camera replacement or strong wind), using outdated scene prior might negatively affect performance. Moreover, Pro3D outputs a 3D detection given a 2D detection. As 2D detections can be false positives, doing so inevitably hallucinates 3D detections which certainly are incorrect. Future work will analyze this aspect in more depth.
@@ 5 Conclusion
  • Section E. We compare the inference speed of Pro3D using different 3D detectors.
  • Section F. We provide our code.
  • Section G. We provide more visual results, including a video demo.
@@ Appendix A Details of Prompt Encoder
Table 10: Results of occluded and truncated obstacles. We split occlusions and truncations with different ranges (0$\sim$50and 50$\sim$100%), respectively. Our approach Pro3D resoundingly outperforms BEVHeight [56] and BEVSpread [48] on all the three superclasses in each range.
(a) Occlusion of 0$\sim$50%
@@
Table 11: Ablation on class-grouping strategies (w.r.t metric AP in BEV). Grouping classes differently with multi-head design (each detector head makes predictions within a specific superclass) makes a difference in training and yields detectors performing differently. Briefly, instead of appearance-based grouping (AG) which groups classes w.r.t object appearance as done by BEVHeight, we do functionality-based grouping (FG), e.g., merging {car, van, truck, bus} into the superclass vehicle. To justify the superiority of FG over AG, we replace the original AG of BEVHeight [56] and BEVSpread [48] with FG, particularly improving on the vehicle class (cf. “BEVHeight w/ FG” vs. “BEVHeight w/ AG”, “BEVSpread w/ FG” vs. “BEVSpread w/ AG”).
@@ Appendix B Loss Design
Class grouping is a common strategy in 3D object detector training. It merges classes into some superclasses, allowing trained features to be shared within superclasses [64]. Moreover, multi-head design follows along with groups of classes for discriminative predictions across fine-grained classes [60]. BEVHeight [56] and BEVSpread [48] adopt class grouping and multi-head design by considering the similarities among object appearance, e.g., size and shape in the 3D world. For example, on DAIR-V2X-I, they create three superclasses that merge {truck, bus}, {car, van}, and {bicyclist, tricyclist, motorcyclist, barrowlist}, respectively, along with the origin pedestrian class. The rationale of such grouping is based on object appearance, e.g., both truck and bus have large size. We call this grouping strategy appearance-based grouping. With such a group strategy, they learn a four-head detector with each making predictions for the corresponding superclass. Differently, we adopt functionality-based grouping, which merges {car, van, truck, bus} into vehicle, and {bicyclist, tricyclist, motorcyclist, barrowlist} into cyclist. This is motivated by the fact that vehicles, cyclists and pedestrians appear at relatively fixed regions on the roadside images. We expect this grouping strategy, or using the location prior of different classes, to facilitate training 3D object detectors on roadside images. Fig. 7 compares the two grouping methods. Perhaps from the same functionality perspective, the DAIR-V2X-I benchmark also merges classes into such superclasses for evaluation. For fair comparison, we use our functionality-based grouping loss to train the 3D detector (e.g., BEVHeight or BEVSpread) within our Pro3D framework to compare with the original 3D detectors. Our extensive study shows that trainig with the functionality-based grouping loss remarkably improves detection performance (Table 11).
@@ Appendix C Analysis of Occlusion and Truncation
Objects might be occluded by others or truncated around the image borders. We carry out breakdown analysis how detectors perform in these scenarios. The DAIR-V2X-I dataset provides occlusion and truncation tags. We split occlusions and truncations with 0$\sim$50and 50$\sim$100for different superclasses on [61], respectively. As shown in Table 10, our method gets better APs on the all superclasses in different ranges of occlusions and truncations. This shows that our method Pro3D is more robust than BEVHeight [56] and BEVSpread [48].
@@ Appendix D Robustness to Noises of Camera Pose
We consider the robustness of Pro3D to noises in camera calibration parameters. We follow BEVHeight [56] to add noises on camera pitch and roll. Table 12 compares the robustness of BEVHeight [56], BEVSpread [48] and our Pro3D. Clearly, Pro3D is more robust than prior works to camera noises.
Table 12: Robustness to camera noise. We follow BEVHeight [56] to simulate calibration errors by adding noises to camera pitch and roll. Results show that Pro3D is more robust than BEVHeight and BEVSPread [48] to these camera pose noises.
@@
Table 13: Comparison of inference time (ms) between different methods. BEVHeight [56] has the fastest inference time comparing with other three methods. Here, we build our Pro3D using it as the 3D detector. It is worth noting that we adopt the functionality-based grouping loss, which induces less parameters compared to the wide-used appearance-based grouping loss (Section B). This makes our Pro3D obtain the fastest inference speed among these methods.
@@
Appendix E Comparison on the Inference Speed
We compare the inference time of different methods in Table 13. BEVHeight [56] gets the fastest inference time in previous work, so we use it in our ablation studies. Moreover, our method Pro3D outperforms other works in inference speed, where we use YOLOV7 as the 2D detector in Pro3D.
Appendix F Open-Source Code
@@
Appendix G More Visualizations and Video Demo
Figure 8 visualizes more results on the DAIR-V2X-I dataset [61]. Pro3D not only has better orientation predictions for vehicles but also can detect hard objects such as infrequently-seen ambulances, occluded pedestrians and cyclists.
Video demo. We include a video demo as a part of the supplementary material, named video-demo.mp4. In this video demo, we compare our Pro3D with the state-of-the-art method BEVSpread [48]. The demo clearly shows that our Pro3D performs much better than BEVSpread in roadside monocular 3D detection.
Figure 8: More visualizations between the state-of-the-art method BEVHeight [56] and our Pro3D.

### v4

@@ 4.4 Societal Impacts and Limitations
Societal impacts. We note potential risks if directly using the proposed Pro3D in real-world applications such as vehicle-vehicle communication and vehicle-infrastructure cooperative perception. These applications should evaluate our approach in their whole system w.r.t safety and robustness. Moreover, as methods (including our Pro3D) developed with existing datasets might be biased towards specific cities and countries. deploying Pro3D requires caution.
Limitations and future work. Currently, our Pro3D exploits a scene prior learned before hand; the change of camera pose (e.g., due to camera replacement or strong wind) can affect performance. While we have studied the robustness of Pro3D to scene priors generated in daytime or at night (Table 6), we did not extend this study to different weather conditions as current datasets did not capture images under diverse weathers. Moreover, Pro3D outputs a 3D detection given a 2D detection. As 2D detections can be false positives, doing so inevitably hallucinates 3D detections which certainly are incorrect. Future work will analyze these aspects in more depth.
@@ 5 Conclusion
  • Section E. We report 2D metrics by roadside monocular 3D detectors.
  • Section F. We conduct cross-dataset evaluation.
  • Section G. We compare the inference speed of Pro3D using different 3D detectors.
  • Section H. We provide our code.
  • Section I. We provide more visual results, including demo videos, and a Jupyter Notebook of code.
@@ Appendix A Details of Prompt Encoder
[…]
@@
Table 11: Ablation on class-grouping strategies (w.r.t metric AP in BEV). Grouping classes differently with multi-head design (each detector head makes predictions within a specific superclass) makes a difference in training and yields detectors performing differently. Briefly, instead of appearance-based grouping (AG) which groups classes w.r.t object appearance as done by BEVHeight, we do functionality-based grouping (FG), e.g., merging {car, van, truck, bus} into the superclass vehicle. To justify the superiority of FG over AG, we replace the original AG of BEVHeight [yang2023BEVHeight] and BEVSpread [wang2024bevspread] with FG, particularly improving on the vehicle class (cf. “BEVHeight w/ FG” vs. “BEVHeight w/ AG”, “BEVSpread w/ FG” vs. “BEVSpread w/ AG”).
@@ Appendix B Loss Design
Class grouping is a common strategy in 3D object detector training. It merges classes into some superclasses, allowing trained features to be shared within superclasses [zhu2019class]. Moreover, multi-head design follows along with groups of classes for discriminative predictions across fine-grained classes [yin2021center]. BEVHeight [yang2023BEVHeight] and BEVSpread [wang2024bevspread] adopt class grouping and multi-head design by considering the similarities among object appearance, e.g., size and shape in the 3D world. For example, on DAIR-V2X-I, they create three superclasses that merge {truck, bus}, {car, van}, and {bicyclist, tricyclist, motorcyclist, barrowlist}, respectively, along with the origin pedestrian class. The rationale of such grouping is based on object appearance, e.g., both truck and bus have large size. We call this grouping strategy appearance-based grouping. With such a group strategy, they learn a four-head detector with each making predictions for the corresponding superclass. Differently, we adopt functionality-based grouping, which merges {car, van, truck, bus} into vehicle, and {bicyclist, tricyclist, motorcyclist, barrowlist} into cyclist. This is motivated by the fact that vehicles, cyclists and pedestrians appear at relatively fixed regions on the roadside images. We expect this grouping strategy, or using the location prior of different classes, to facilitate training 3D object detectors on roadside images. Fig. 8 compares the two grouping methods. Perhaps from the same functionality perspective, the DAIR-V2X-I benchmark also merges classes into such superclasses for evaluation. For fair comparison, we use our functionality-based grouping loss to train the 3D detector (e.g., BEVHeight or BEVSpread) within our Pro3D framework to compare with the original 3D detectors. Our extensive study shows that trainig with the functionality-based grouping loss remarkably improves detection performance (Table 11).
@@ Appendix C Analysis of Occlusion and Truncation
[…]
@@ Appendix D Robustness to Noises of Camera Pose
We consider the robustness of Pro3D to noises in camera calibration parameters. We follow BEVHeight [yang2023BEVHeight] to add noises on camera pitch and roll. Table 12 compares the robustness of BEVHeight [yang2023BEVHeight], BEVSpread [wang2024bevspread] and our Pro3D. Clearly, Pro3D is more robust than prior works to camera noises.
Table 12: Robustness to camera noise. We follow BEVHeight [yang2023BEVHeight] to simulate calibration errors by adding noises to camera pitch and roll. Results show that Pro3D is more robust than BEVHeight and BEVSPread [wang2024bevspread] to these camera pose noises.
@@
Table 13: Comparison of inference time (ms) between different methods. BEVHeight [yang2023BEVHeight] has the fastest inference time comparing with other three methods. Here, we build our Pro3D using it as the 3D detector. It is worth noting that we adopt the functionality-based grouping loss, which induces less parameters compared to the wide-used appearance-based grouping loss (Section B). This makes our Pro3D obtain the fastest inference speed among these methods.
@@
Appendix E 2D Detection Performance of 3D Detectors
Table 14 reports 2D metrics of both the 2D detector DINO and roadside 3D detection methods including BEVDepth, BEVHeight, BEVSpread, and their improved versions by our Pro3D. Results demonstrate that (1) existing 3D detectors (BEVDepth, BEVHeight, and BEVSpread) produce significantly lower 2D metrics than the 2D detector DINO; (2) by exploiting DINO’s 2D detections as prompt, our Pro3D greatly enhances 3D detection (see results in Table 1 and 2 in the main paper).
Table 14: Comparison w.r.t 2D detection metric (mAP) on DAIR-V2X-I. For both 2D detector (DINO) and 3D detectors BEVDepth and BEVHeight, we train them on the training set of DARI-V2X-I. To evaluate 3D detectors w.r.t 2D metrics, we project their 3D detections onto the 2D image plane to derive their predicted 2D boxes. Compared to BEVDepth and BEVHeight, DINO achieves >>10 mAP higher on all classes and $\sim$20 mAP higher on Pedestrian! Owing to the effective leverage of the 2D detector DINO, our Pro3D significantly boosts performance of the compared methods.
Method Vehicle Cyclist Pedestrian
Easy Mid Hard Easy Mid Hard Easy Mid Hard
2D detector (DINO) [zhang2022dino] 74.3 75.6 75.6 57.3 54.7 54.9 50.6 51.6 51.6
BEVDepth [li2023BEVDepth] 62.9 64.8 64.2 41.5 44.8 44.9 32.9 33.1 33.0
Pro3D w/ BEVDepth 68.0 68.3 68.9 45.9 47.7 47.6 42.1 42.6 42.6
BEVHeight [yang2023BEVHeight] 63.5 64.7 64.5 41.4 44.6 44.7 33.6 33.7 33.9
Pro3D w/ BEVHeight 68.1 68.8 68.8 47.9 49.5 49.4 42.8 43.5 43.5
BEVSpread [wang2024bevspread] 64.1 65.2 64.6 41.7 45.1 44.9 33.9 34.1 34.3
Pro3D w/ BEVSpread 68.7 69.4 69.2 48.7 50.3 50.3 44.1 44.9 44.9
Appendix F Cross-Dataset Evaluation
One may wonder the results of cross-dataset evaluation. The datasets (DAIR-V2X-I and Rope3D) used in our work do not have the same vocabulary, but both have the vehicle class. Therefore, focusing on the vehicle class, we use the models trained on Rope3D and test on the DAIR-V2X-I benchmark. In this study, we compare our Pro3D that uses the DINO 2D-detector. The table below compares BEVHeight, BEVSpread, and Pro3D. Performance of all methods degrade (marked in parentheses) but our Pro3D yields less degradation compared with the other methods.
Method on vehicle Easy Mid Hard
BEVHeight [yang2023BEVHeight] 62.6 (-15.2) 50.1 (-15.7) 50.3 (-15.6)
Pro3D w/ BEVHeight 70.0 (-10.7) 60.7 (-11.1) 61.1 (-10.9)
BEVSpread [wang2024bevspread] 65.9 (-13.2) 53.0 (-13.8) 53.3 (-13.6)
Pro3D w/ BEVSpread 77.6 (-7.6) 65.1 (-8.0) 65.3 (-8.3)
Appendix G Comparison on the Inference Speed
We compare the inference time of different methods in Table 13. BEVHeight [yang2023BEVHeight] gets the fastest inference time in previous work, so we use it in our ablation studies. Moreover, our method Pro3D outperforms other works in inference speed, where we use YOLOV7 as the 2D detector in Pro3D.
Appendix H Open-Source Code
@@
Appendix I More Visualizations, Demo Videos, and Jupyter Notebook of Code
Figure 9 visualizes more results on the DAIR-V2X-I dataset [yu2022dair]. Pro3D not only has better orientation predictions for vehicles but also can detect hard objects such as infrequently-seen ambulances, occluded pedestrians and cyclists.
Demo Videos. We include two demo videos as a part of the supplementary material, named demo-video-vs-BEVHeight.mp4 and demo-video-vs-BEVSpread.mp4. These videos compare our Pro3D with BEVHeight [yang2023BEVHeight] and BEVSpread [wang2024bevspread], respectively. They clearly show that our Pro3D performs much better than both methods in roadside monocular 3D detection.
Jupyter Notebook of Code. In the supplement, we provide not only our code but also a Jupyter Notebook file named demo-pro3d-infer-vis.ipynb which can readily run our Pro3D and display its results on video frame examples. The reader is referred to this file for a quick visualization and guideline on how to run our code.
Figure 9: More visualizations between the state-of-the-art method BEVHeight [yang2023BEVHeight] and our Pro3D.